The master budget at Western Company last period called for sales of 225,800 units at $9.8 each. The costs were estimated to be $3.83 variable per unit and $225,800 fixed. During the period, actual production and actual sales were 230,800 unit. The selling price was $9.90 per unit. Variable costs were $5.30 per unit. Actual fixed costs were $225,800.

Required:

Prepare a sales activity variance analysis. (Indicate the effect of each variance by selecting “F” for favorable, or “U” for unfavorable. If there is no effect, do not select either option.)

WESTERN COMPANY

Sales Activity Variance

Flexible Budget

Sales Activity Variance

Master Budget

Sales revenue

Less:

   Variable costs

Contribution margin

Less:

   Fixed costs

Operating profits
Answer #1
WESTERN COMPANY
Sales Activity Variance
Flexible Budget Sales Activity Variance Master Budget
Sales revenue 2261840 49000 F 2212840
Less:
Variable costs 883964 19150 U 864814
Contribution margin 1377876 29850 F 1348026
Less:
Fixed costs 225800 0 225800
Operating profits 1152076 29850 F 1122226
Workings:
Flexible Budget Master Budget
Sales revenue =230800*9.80 =225800*9.80
Variable costs =230800*3.83 =225800*3.83